OpenStudy (anonymous):

How many combinations of 5 students can a teacher choose from 24 students? A. 120 B. 42,504 C. 5,100,480 D. 7,962,624

OpenStudy (anonymous):

i would perfer righting this out

OpenStudy (nirmalnema):

24C5 24!/(24-5)!

OpenStudy (bradely):

yes nirmalnena answer correct

OpenStudy (nirmalnema):

thanks @bradely

OpenStudy (bradely):

but small modification

OpenStudy (bradely):

24C5 =24!/(24-5)!5!

OpenStudy (anonymous):

whats the answer?

OpenStudy (nirmalnema):

24*23*22*21*20*19! / 19! 5! =24*23*22*21*20 /5!

OpenStudy (bradely):

24*23*22*21*20*19!/(19!)(5*4*3*2*1)=24*23*22*21*20*/)(5*43**2*1 =42504

OpenStudy (nirmalnema):

=42504

OpenStudy (nirmalnema):

@bradely is correct

OpenStudy (anonymous):

thank u guys for ur time and hard work

OpenStudy (bradely):

thank u nirmalnena